• 21:16 UT: lost scans due to SpaceX tracking test. schedule halted after scan 291-2112 at 21:13:28. Scheduled resumed with scan 291-2153 at 21:53:01. Lost scans: 291-2114-291-2151. (Lucia)
  • strong RFI in last two channels (Lucia).
  • Recording to 8TB module HOB+0088 in bank B, +0GB (Ross).
  • 23:50 UT: Antenna drifting off and subsequently re-acquiring source, Az: -124.77 El: 29.84 (Ross).
  • 00:00 UT: Experiment started OK (Ross).


  • s-band temperature high, two bad channels due DBBC problems - this is a known problem (Lucia).
  • Recording to 16TB module HOB+0108 in bank A, +126GB (Ross).
  • 00:00 UT: Experiment started OK (Ross).


  • Recording to 16TB module HOB+0100 in bank A, +0GB (Ross).
  • 22:30 UT: Network connection down, switched over from ADSL to 3G (Ross).
  • 00:00 UT: Experiment started OK (Ross).